# Sparse dynamic network reconstruction through $L_{1}$-regularization of a Lyapunov equation

### v2

@@ Abstract
An important problem in many areas of science is that of recovering interaction networks from simultaneous time-series of many interacting dynamical processes. A common approach is to use the elements of the correlation matrix or its inverse as proxies of the interaction strengths, but the reconstructed networks are necessarily undirected. Transfer entropy methods have been proposed to reconstruct directed networks but the reconstructed network lacks information about interaction strengths. We propose a network reconstruction method that inherits the best of the two approaches by reconstructing a directed weighted network from noisy data under the assumption that the network is sparse and the dynamics are governed by a linear (or weakly-nonlinear) stochastic dynamical system. The two steps of our method are i) constructing an (infinite) family of candidate networks by solving the covariance matrix Lyapunov equation for the state matrix and ii) using $L_{1}$-regularization to select a sparse solution. We further show how to use prior information on the (non)existence of a few directed edges to drastically improve the quality of the reconstruction.
{IEEEkeywords}
network identification, network inference, stochastic network dynamics, linear programming
1 Introduction
In many areas of science it is common to record simultaneous time-series of interacting dynamical processes. A natural and important problem is then to develop data-driven methods that can recover, or reconstruct, the underlying interaction network, possibly including information about both the direction and the strength of the interactions. Reconstructing interaction networks has applications, e.g., in neuroscience (between neurons or neuronal populations), in ecology (i.e., between species populations in a community), in molecular biology (i.e., between expression of genes), and in social sciences (i.e., between opinionated individuals).
@@ 1 Introduction
We introduce a novel method to reconstruct directed and weighted dynamical networks under the assumption that the network dynamics can be modelled by a linear (or weakly non-linear) stochastic differential equation driven by a white noise vector. When used to describe resting state brain activity, this model is known as the linear Dynamical Causal Model [10, 11]. Under this assumption, the covariance matrix of the model is the solution to a Lyapunov equation that links it to the state matrix of the network [12, 13]. If the state matrix is assumed to be symmetric (i.e., the network is undirected) and the noise components are uncorrelated, then solving the Lyapunov equation for the state matrix returns exactly the precision matrix [1]. But if no symmetry assumptions are made, then solving the Lyapunov equation for the state matrix returns a whole affine subspace of candidate solutions. Under the extra assumption that the interaction network is sparse, $L_{1}$-regularization can effectively select a sparse solution from among the candidates. Prior knowledge on edge existence can be included in the constraints of the resulting linear programming (LP) problem to enhance the method performance.
@@
The main contribution of the paper is to formulate an $L_{1}$-regularization problem for sparse network reconstruction as an LP optimization problem over the affine solution space of all state matrices that solve the covariance Lyapunov equation for a given covariance matrix. Crucially, this requires the preliminary construction of an isomorphism from such a solution space to the solution space of an underdetermined linear system of equations. The second contribution is to show how to introduce prior knowledge on edge existence into the constraints of the LP problem. The third contribution is to suggest the use of a TE-based network inference algorithm as a way of obtaining some priors on edge existence and show that it significantly improves the quality of network reconstruction. As an auxiliary contribution, we propose a simple algorithm to construct sparse Hurwitz matrices.
@@
The paper is structured as follows. Notation is introduced in Section 2. The network model and its interpretation are introduced in Section 3. Section 4 studies the geometry of the solution space of the covariance Lyapunov equation for the state matrix and introduces the isomorphism needed for the LP formulation of an $L_{1}$-regularization problem over this space. Section 5 explicitly describes the LP formulation of the $L_{1}$-regularization problem. Section 6 shows how prior information about edge existence can be incorporated into the constraints of the LP problem. Section 7 introduces an algorithm based on TE to provide prior knowledge on edge existence and show how to use it to constrain the LP problem. Finally, Section 8 presents extensive numerical validation results obtained over large families of randomly generated sparse Hurwitz matrices. Conclusions and perspectives are presented in Section 9.
2 Notation
@@
We denote $[n]=\{1,...,n\}$. Given $n,m\in\mathbb{N}$, we write $\mathbb{R}^{m\times n}$ for the set of $m$ by $n$ matrices with real entries. For $i\in[m]$, and $j\in[n]$, the value at the $i$-th row and $j$-th column of $A\in\mathbb{R}^{m\times n}$ is $A_{i,j}$. The column-stack vectorization isomorphism $\text{vec}:\mathbb{R}^{m\times n}\to\mathbb{R}^{mn}$ is given by $(\text{vec}(A))_{m(j-1)+i}=A_{i,j}$. Vector $\mathbf{1}_{n}\in\mathbb{R}^{n}$ consists of $n$ entries equal to $1$ (and similarly for $\mathbf{0}_{n}$). The set of eigenvalues of $B\in\mathbb{R}^{n\times n}$ is $\sigma(B)$. The real part of $z\in\mathbb{C}$ is denoted $\Re(z)$.
3 Model formulation
@@
We consider a network model described by the stochastic differential equation (SDE) [13]
@@
where $w$ is a white noise vector (with spectral density matrix $Q=I$) and $A\in\mathbb{R}^{n\times n}$ is a Hurwitz matrix, i.e., all of its eigenvalues have a strictly negative real part. Any solution $x$ for system (1) is a stochastic process with mean 00 and covariance matrix $\Gamma=\mathbf{E}(xx^{T})$ (see for example Section $4.3$ in[13]).
@@
4 Solving the covariance Lyapunov equation for the state matrix
Figure 1: Heatmaps for three examples of weighted connectivity matrices of dimension $8\times 8$. The leftmost heatmap is for a sparse matrix $A\in\mathbb{R}^{8\times 8}$ generated randomly using the algorithm in Section 8. The other two matrices were selected randomly from the solution space $\mathcal{S}_{\Gamma}$ for the matrix $A$. This means the solution to the system in (1) has the same covariance matrix for all three matrices.
@@ 4 Solving the covariance Lyapunov equation for the state matrix
Given $A$, the stable covariance matrix solution can be computed by
$\text{vec}(\Gamma)=-(I\otimes A+A\otimes I)^{-1}\text{vec}(I).$
However, because the covariance matrix is symmetric, the inverse problem, i.e., solving for $A$ given $\Gamma$, does not have a unique solution [14]. Following [14], given $\Gamma$, define its orthonormal spectral decomposition:
@@
The set of matrices $\bar{A}$ satisfying (4) is an affine subspace of dimension $m=\frac{n(n-1)}{2}$. Since $A=U\bar{A}U^{T}$, this means that the set of possible solutions $A$ to the inverse covariance Lyapunov equations is also an $m$-dimensional affine space $\mathcal{S}_{\Gamma}\subseteq\mathbf{R}^{n\times n}$. Figure 1 shows a Hurwitz matrix $A$, with correlation matrix $\Gamma$, together with other two matrices taken randomly from the solution space $\mathcal{S}_{\Gamma}$, hence associated to the same covariance matrix.
@@
$\mathbf{M}_{(i,j)(\ell,k)}=\begin{cases}c_{j}U_{\ell,i}U_{k,j}+c_{i}U_{\ell,j}U_{k,i}&\text{if }i\neq j\\ U_{\ell,i}U_{k,j}&\text{if }i=j\end{cases}$
@@
$\mathbf{b}_{(i,j)}=\begin{cases}0&\text{if }i\neq j\\ \frac{-1}{2c_{i}}&\text{if }i=j\end{cases}$
[…]
Proof 4.2.
[…]
@@
[…]
5 $L_{1}$-regularization on $\mathcal{S}_{\Gamma}$ and sparse network reconstruction
@@
In turn, (7) can be transformed into a Linear Programming (LP) problem by taking a new variable $u\in\mathbb{R}^{n^{2}}$ and a vector $c=(\mathbf{1}_{n^{2}},\mathbf{0}_{n^{2}})$, and defining the LP problem:
Minimize $c^{T}$ $\begin{pmatrix}u\\ v\end{pmatrix}$ (10)
Subject to $\begin{bmatrix}\mathbf{0}&\mathbf{M}\end{bmatrix}$ $\begin{pmatrix}u\\ v\end{pmatrix}=\mathbf{b}$ (14)
and $\begin{bmatrix}-I&I\\ -I&-I\end{bmatrix}$ $\begin{pmatrix}u\\ v\end{pmatrix}\leq\mathbf{0}_{2n^{2}}.$ (19)
The LP problem (10) aims at minimizing the objective $\sum_{(i,j)}u_{(i,j)}$ subject to the restriction that $\text{vec}^{-1}(v)\in\mathcal{S}_{\Gamma}$ and $\lvert v_{(i,j)}\rvert\leq u_{(i,j)}$ for all $i,j\in[n]$. A solution to the LP problem (10) immediately gives a solution to the problem (7) and vice-versa – in this sense the two problems are equivalent.
Solving (10) will return a sparse solution from $\mathcal{S}_{\Gamma}$. However, because there exist multiple sparse solutions in $\mathcal{S}_{\Gamma}$, the returned sparse solution might still be very different from the true network adjacency matrix, mostly depending on the initial problem data. In the following section we present a way to constrain the optimization problem by including prior approximate knowledge on edge existence to bias the reconstruction toward solutions that are consistent with the priors.
6 Including priors on edge existence
@@
The idea for including priors comes from the observation that if we knew that exactly $p$ directed edges (and which ones) existed in the network, then $A$ would lie in $\mathcal{R}_{\Gamma}\subseteq\mathbf{R}^{n\times n}$ of dimension $p$, i.e., the subspace parameterized by the $p$ nonzero elements of $A$. By Proposition 4.1, $A\in\mathcal{S}_{\Gamma}$. Thus, $A\in\mathcal{S}_{\Gamma}\cap\mathcal{R}_{\Gamma}$, that is, $A$ must lie in the intersection of two affine subspaces of dimension $m=\dim(\mathcal{S}_{\Gamma})=\frac{n(n-1)}{2}$ and $p=\dim(\mathcal{R}_{\Gamma})$, respectively.
It is well known (a consequence of Theorem 2.43 in[16]) that the transversal intersection of two linear subspaces of $\mathbb{R}^{n\times n}$ of dimension $p$ and $m$, with $m+p\leq n^{2}$ is either empty or a point. This means that if we have exact knowledge of which edges exist in the network and the network is sparse enough (i.e., $m+p\leq n^{2}$), then solving (10) restricted to $\mathcal{R}_{\Gamma}$ would return $A$ exactly. However, given only imperfect knowledge of edge existence and if $m+p<n^{2}$ (which is the case for moderately sparse networks), then our noisy approximation ^$\hat{\mathcal{R}}_{\Gamma}$ of $\mathcal{R}_{\Gamma}$ and $\mathcal{S}_{\Gamma}$ will almost surely not intersect at all. In spite of this, looking for the matrix $A\in\mathcal{S}_{\Gamma}$ with minimal distance to ^$\hat{\mathcal{R}}_{\Gamma}$ should bias the solution (10) toward the real $A$. This is the approach we use here.
Based on this idea we define a vector $Z\in[0,1]^{n^{2}}$ with entries representing the relative importance assigned to each connection during the optimization. The objective of the minimization then becomes $\sum_{(i,j)}Z_{(i,j)}\lvert v_{(i,j)}\rvert$. In the simplest case, we can introduce information about the presence of edges in the following way. Let =$\mathcal{I}=\{(i,j)\in[n]^{2}\mid(i,j)\text{ is a known edge}\}$. Then we can define $Z$ as
@@
Minimize $c^{T}$ $\begin{pmatrix}u\\ v\end{pmatrix}$ (22)
Subject to $\begin{bmatrix}\mathbf{0}&\mathbf{M}\end{bmatrix}$ $\begin{pmatrix}u\\ v\end{pmatrix}=\mathbf{b}$ (26)
and $\begin{bmatrix}-I&\text{diag}(Z)\\ -I&-\text{diag}(Z)\end{bmatrix}$ $\begin{pmatrix}u\\ v\end{pmatrix}\leq\mathbf{0}_{2n^{2}}.$ (31)
@@
7 Edge priors from Transfer Entropy
There are various ways to obtain prior knowledge on edge existence, e.g., from expert knowledge or from previous analysis. One generally applicable tool for inferring the existence of directed interactions is Transfer Entropy (TE)[4]. Given two discrete-time stationary stochastic processes $X$ and $Y$, the TE at $T_{X\to Y}$ from $X$ to $Y$ is a measure of the predictive power of $X$ about $Y$. In its simplest form $T_{X\to Y}({\tt t})$, the transfer entropy $T_{X\to Y}$ at (discrete) time ${\tt t}$, can be defined as [4]

### v3

@@ Abstract
An important problem in many areas of science is that of recovering interaction networks from high-dimensional time-series of many interacting dynamical processes. A common approach is to use the elements of the correlation matrix or its inverse as proxies of the interaction strengths, but the reconstructed networks are necessarily undirected. Transfer entropy methods have been proposed to reconstruct directed networks, but the reconstructed network lacks information about interaction strengths. We propose a network reconstruction method that inherits the best of the two approaches by reconstructing a directed weighted network from noisy data under the assumption that the network is sparse and the dynamics are governed by a linear (or weakly-nonlinear) stochastic dynamical system. The two steps of our method are i) constructing an (infinite) family of candidate networks by solving the covariance matrix Lyapunov equation for the state matrix and ii) using $L_{1}$-regularization to select a sparse solution. We further show how to use prior information on the (non)existence of a few directed edges to dramatically improve the quality of the reconstruction.
I Introduction
In many areas of science it is common to record the evolution in time of many interacting processes. A natural and important problem is then to develop data-driven methods that can recover, or reconstruct, from the resulting high-dimensional time-series the underlying interaction network. Ideally, the reconstructed network would include information about both the direction and the strength of the interactions. Interaction networks reconstruction has applications, e.g., in neuroscience (between neurons or neuronal populations), in ecology (i.e., between species populations in a community), in molecular biology (i.e., between expression of genes), and in social sciences (i.e., between opinionated individuals).
@@ I Introduction
A commonly used, mathematically tractable approximation is to assume that the network dynamics can be modelled by a linear (or weakly nonlinear) stochastic differential equation driven by a white noise vector. When used to describe resting-state brain activity, the resulting model is known as the (linear) Dynamical Causal Model (DCM) [10, 11]. Under this assumption, the covariance matrix of the model is the solution to a Lyapunov equation that links it to the state matrix of the network [12, 13]. If the state matrix is assumed to be symmetric (i.e., the network is undirected) and the noise components are uncorrelated, then solving the Lyapunov equation for the state matrix returns exactly the precision matrix [1]. But if no symmetry assumptions are made, then solving the Lyapunov equation for the state matrix returns a whole affine subspace of candidate solutions. Methods using differential covariance [2] enable singling out asymmetric solutions, but these methods require computing the time-derivative of the measured time-series, which can problematic or unfeasible when measurement or intrinsic noise are high.
We introduce a novel method to reconstruct weighted, thus directed, dynamical networks, using the idea that under the extra assumption that the interaction network is sparse, $L_{1}$-regularization can effectively select such a sparse solution from among the candidates. We also show how prior knowledge of edge existence can be included in the constraints of the resulting linear programming (LP) problem to enhance the method performance.
@@
The first and main contribution of the paper is to formulate an $L_{1}$-regularization problem for sparse network reconstruction as an LP optimization problem over the affine solution space of all state matrices that solve the covariance Lyapunov equation for a given covariance matrix. Crucially, this requires the preliminary construction of an isomorphism from such a solution space to the solution space of an underdetermined linear system of equations. The second contribution is to show how to introduce prior knowledge on edge existence into the constraints of the LP problem. The third contribution is to suggest the use of a TE-based network inference algorithm as a way of obtaining some priors on edge existence and show that it significantly improves the quality of network reconstruction.
@@
The paper is structured as follows. Notation is introduced in Section II. The network model and its interpretation are introduced in Section III. Section IV studies the geometry of the solution space of the covariance Lyapunov equation for the state matrix and introduces the isomorphism needed for the LP formulation of an $L_{1}$-regularization problem over this space. Section V explicitly describes the LP formulation of the $L_{1}$-regularization problem. Section VI shows how prior information about edge existence can be incorporated into the constraints of the LP problem. Section VII introduces an algorithm based on TE to provide prior knowledge on edge existence and show how to use it to constrain the LP problem. Finally, Section VIII presents extensive numerical validation results obtained over large families of randomly generated sparse Hurwitz matrices. Conclusions and perspectives are presented in Section IX.
II Notation
@@
We denote $n]=\{1,...,n\}$. Given $n,m\in\mathbb{N}$, we write $\mathbb{R}^{m\times n}$ for the set of $m$ by $n$ matrices with real entries. For $i\in[m]$, and $j\in[n]$, the value at the $i$-th row and $j$-th column of $A\in\mathbb{R}^{m\times n}$ is $A_{i,j}$. The column-stack vectorization isomorphism $\text{vec}:\mathbb{R}^{m\times n}\to\mathbb{R}^{mn}$ is given by $(\text{vec}(A))_{m(j-1)+i}=A_{i,j}$. Vector $\mathbf{1}_{n}\in\mathbb{R}^{n}$ consists of $n$ entries equal to $1$ (and similarly for $\mathbf{0}_{n}$). The set of eigenvalues of $B\in\mathbb{R}^{n\times n}$ is $\sigma(B)$. The real part of $z\in\mathbb{C}$ is denoted $\Re(z)$. For a pair of matrices $A\in\mathbb{R}^{n\times n},B\in\mathbb{R}^{n\times n}$, $\otimes$ is the Kronecker product given by $\forall r,v,s,w\in[n],(A\otimes B)_{n(r-1)+v,n(s-1)+w}=(A_{r,s})(B_{v,w})$, so that $A\otimes B\in\mathbb{R}^{n^{2}\times n^{2}}$.
III Model formulation
@@
We consider a network model described by the linear stochastic differential equation (SDE) [13]
@@
where $w$ is a white noise vector (with noise covariance $Q=I$) and $A\in\mathbb{R}^{n\times n}$ is a Hurwitz matrix, i.e., all of its eigenvalues have a strictly negative real part. Any solution $x$ for system (1) is a stochastic process with mean 0 and covariance matrix $\Gamma=\mathbf{E}(xx^{T})$ (see for example Section $4.3$ in[13]).
@@
IV Solving the covariance Lyapunov equation for the state matrix
Figure 1: Heatmaps for three examples of weighted connectivity matrices of dimension $8\times 8$. The leftmost heatmap is for a sparse matrix $A\in\mathbb{R}^{8\times 8}$ generated randomly using the algorithm in Section VIII. The other two matrices were selected randomly from the solution space $\mathcal{S}_{\Gamma}$ for the matrix AA. This means the solution to the system in (1) has the same covariance matrix for all three matrices.
@@ IV Solving the covariance Lyapunov equation for the state matrix
Given $A$, the stable covariance matrix solution is given by
$\text{vec}(\Gamma)=-(I\otimes A+A\otimes I)^{-1}\text{vec}(I),$
which is readily computed. However, because the covariance matrix is symmetric, the inverse problem, i.e., solving for $A$ given $\Gamma$, does not have a unique solution [14]. The differential covariance approach [2] uses the fact that the expected value of $\dot{x}x^{T}$ is equal to $A\Gamma$ to obtain a unique inverse solution $A$. However, when measurement or intrinsic noise are large, computing time-series derivatives can be problematic or unfeasible. Our goal is to propose a new algorithmic method that, under suitable assumptions, can single out a unique solution to the inverse covariance Lyapunov equation (2). Following [14], given $\Gamma$, define its orthonormal spectral decomposition:
@@
The set of matrices $\bar{A}$ satisfying (4) is an affine subspace of dimension $m=\frac{n(n-1)}{2}$. Since $A=U\bar{A}U^{T}$, this means that the set of possible solutions $A$ to the inverse covariance Lyapunov equations is also an $m$-dimensional affine space $\mathcal{S}_{\Gamma}\subseteq\mathbb{R}^{n\times n}$. Figure 1 shows a Hurwitz matrix $A$, with correlation matrix $\Gamma$, together with other two matrices taken randomly from the solution space $\mathcal{S}_{\Gamma}$, hence associated to the same covariance matrix.
@@
$\mathbf{M}_{\iota_{1}(i,j),\iota_{2}(\ell,k)}=\begin{cases}c_{j}U_{\ell,i}U_{k,j}+c_{i}U_{\ell,j}U_{k,i}&\text{if }i\neq j\\ U_{\ell,i}U_{k,j}&\text{if }i=j\end{cases}$
@@
$\mathbf{b}_{\iota_{1}(i,j)}=\begin{cases}0&\text{if }i\neq j\\ \frac{-1}{2c_{i}}&\text{if }i=j\end{cases}$
where $\iota_{1}:\{(i,j)\mid i,j\in[n],i\leq j\}\to[\frac{n(n+1)}{2}]$ and $\iota_{2}:[n]^{2}\to[n^{2}]$ are one-to-one functions that give the index of the vectorizations of the upper triangle of $\bar{A}$ (including the diagonal) and of the full matrix respectively. Then for $v\in\mathbb{R}^{n^{2}}$, we have that $\text{vec}^{-1}(v)\in\mathcal{S}_{\Gamma}$ if and only if $\mathbf{M}v=\mathbf{b}$.
Proof:
Let $v\in\mathbb{R}^{n^{2}}$ and take $A=\text{vec}^{-1}(v)$. Suppose $A\in\mathcal{S}_{\Gamma}$. From (4), if $i\neq j$ then $\sum_{\ell=1}^{n}\sum_{k=1}^{n}U_{\ell,i}A_{\ell,k}U_{k,j}=(U^{T}AU)_{i,j}=\bar{A}_{i,j}=-\frac{c_{i}}{c_{j}}\bar{A}_{j,i}=-\frac{c_{i}}{c_{j}}\sum_{\ell=1}^{n}\sum_{k=1}^{n}U_{\ell,j}A_{\ell,k}U_{k,i}$, from which it follows that
@@
Following the same argument, for any $v\in\mathbb{R}^{n^{2}}$, if $\mathbf{M}v=\mathbf{b}$, then $\text{vec}^{-1}(v)\in\mathcal{S}_{\Gamma}$. ∎
V $L_{1}$-regularization on $\mathcal{S}_{\Gamma}$ and sparse network reconstruction
@@
In turn, (7) can be transformed into a Linear Programming (LP) problem by taking a new variable $u\in\mathbb{R}^{n^{2}}$ and a vector $\mathfrak{s}=(\mathbf{1}_{n^{2}},\mathbf{0}_{n^{2}})$, and defining the LP problem:
Minimize ${\mathfrak{{s}}}^{T}$ $\begin{bmatrix}u\\ v\end{bmatrix}$ (8)
Subject to $\begin{bmatrix}\mathbf{0}&\mathbf{M}\end{bmatrix}$ $\begin{bmatrix}u\\ v\end{bmatrix}=\mathbf{b}$
and $\begin{bmatrix}-I&I\\ -I&-I\end{bmatrix}$ $\begin{bmatrix}u\\ v\end{bmatrix}\leq\mathbf{0}_{2n^{2}}.$
The LP problem (8) minimizes the objective $\sum_{(i,j)}u_{(i,j)}$ subject to the restriction that $\text{vec}^{-1}(v)\in\mathcal{S}_{\Gamma}$ and $\lvert v_{(i,j)}\rvert\leq u_{(i,j)}$ for all $i,j\in[n]$. A solution to the LP problem (8) immediately gives a solution to the problem (7) and vice-versa – in this sense the two problems are equivalent [15].
Solving (8) will return a sparse solution from $\mathcal{S}_{\Gamma}$. However, because there might be (and usually are) many sparse matrices in $\mathcal{S}_{\Gamma}$, the solution to the $L_{1}$-regularization problem might still be very different from the true network adjacency matrix, including under the sparsity assumption and mostly depending on the initial problem data. In the following section we present a way to constrain the optimization problem by including approximate prior knowledge on edge existence to bias the reconstruction toward solutions that are consistent with the priors.
VI Including priors on edge existence
@@
The idea for including priors comes from the observation that if we knew that exactly $p$ directed edges (and which ones) existed in the network, then $A$ would lie in $\mathcal{R}_{\Gamma}\subseteq\mathbf{R}^{n\times n}$ of dimension $p$, i.e., the subspace parameterized by the $p$ nonzero elements of $A$. By Proposition IV.1, $A\in\mathcal{S}_{\Gamma}$. Thus, $A\in\mathcal{S}_{\Gamma}\cap\mathcal{R}_{\Gamma}$, that is, $A$ must lie in the intersection of two affine subspaces of dimension $m=\dim(\mathcal{S}_{\Gamma})=\frac{n(n-1)}{2}$ and $p=\dim(\mathcal{R}_{\Gamma})$, respectively.
It is well known (a consequence of Theorem 2.43 in[16]) that the transversal intersection of two linear subspaces of $\mathbb{R}^{n\times n}$ of dimension $p$ and $m$, with $m+p\leq n^{2}$ is either empty or a point. This means that if we have exact knowledge of which edges exist in the network and the network has a sufficiently small number of connections (i.e., $m+p\leq n^{2}$), then solving (8) restricted to $\mathcal{R}_{\Gamma}$ would return $A$ exactly. However, given only imperfect knowledge of edge existence and if $m+p\leq n^{2}$, then our noisy approximation $\hat{\mathcal{R}}_{\Gamma}$ of $\mathcal{R}_{\Gamma}$ and $\mathcal{S}_{\Gamma}$ will almost surely not intersect at all. In spite of this, looking for the matrix $A\in\mathcal{S}_{\Gamma}$ with minimal distance to $\hat{\mathcal{R}}_{\Gamma}$ should bias the solution (8) toward the real $A$. This is the approach we use here.
Based on this idea we define a vector with entries representing the relative importance assigned to each connection during the optimization, $Z\in[0,1]^{n^{2}}$. The objective of the minimization then becomes $\sum_{(i,j)}Z_{(i,j)}\lvert v_{(i,j)}\rvert$. In the simplest case, we can introduce information about the presence of edges in the following way. Let $\mathcal{I}=\{(i,j)\in[n]^{2}\mid(i,j)\text{ is a known edge}\}$. Then we can define $Z$ as
@@
Minimize ${\mathfrak{s}}^{T}$ $\begin{bmatrix}u\\ v\end{bmatrix}$ (9)
Subject to $\begin{bmatrix}\mathbf{0}&\mathbf{M}\end{bmatrix}$ $\begin{bmatrix}u\\ v\end{bmatrix}=\mathbf{b}$
and $\begin{bmatrix}-I&\text{diag}(Z)\\ -I&-\text{diag}(Z)\end{bmatrix}$ $\begin{bmatrix}u\\ v\end{bmatrix}\leq\mathbf{0}_{2n^{2}}.$
@@
VII Edge priors from Transfer Entropy
There are various ways to obtain prior knowledge on edge existence, e.g., from expert knowledge or from previous analysis. One generally applicable tool for inferring the existence of directed interactions is Transfer Entropy (TE)[4]. Given two discrete-time stationary stochastic processes $X$ and $Y$, the TE from $X$ to $Y$ is a measure of the predictive power of $X$ about $Y$ at a discrete time $\tt{t}$. In its simplest form TE is defined as [4]